what is the base area and height of a rectangle whose perimeter is 44in and the base is two less than three times the height.
Answer by ewatrrr(24785)


Hi,
Let x and (3x-2in) represent the height and base respectively,
Question states***
Perimeter = 44in = x + x +(3x-2in) + (3x-2in)
44in = x + x +(3x-2in) + (3x-2in)
solving for x
44in = 8x - 4in
48in = 8x
6 = x, 6in is the height, the base is 16in (18in-2in)
CHECKING our Answer
2*6in + 2*16in = 12in + 32in = 44in